Graph the first-degree inequality in two unknowns. with points
y > 4x − 4

Answer by llenomcdaryl(4)

y > 4x − 4
you will just change the sign. > to =
y=4x-4
Let x= 0
y=4x-4
y=4(0)-4
y=0-4
y=-4
First point (0,-4)
Let x= -1
y=4x-4
y=4(-1)-4
y=-4-4
y=0
Second point (-1,0)
Graph the points in a plane. < or > broken line >= or <= solid line
Substitute (0,0) (recommended) to the original inequality
y > 4x − 4
0 > 4(0)-4
0 > 0-4
0 > -4
If true, shade the part where (0,0) belongs
If false, shade the opposite part where (0,0) doesn't belong